# Representation Consistency for Accurate and Coherent LLM Answer Aggregation

### v1

@@ 2 Problem Definition
A QA task is a tuple $\tau=(\mathcal{Q},\mathcal{A},\rho,phrase,parse,\ell)$. $\mathcal{Q}$ is a set of questions and $\mathcal{A}$ is a set of answers. $\rho\in\Delta(\mathcal{Q}\times\mathcal{A})$ is a distribution from which questions $q$ and ground truth answers $a^{*}$ are sampled. $phrase:\mathcal{Q}\rightarrow\Delta(\mathcal{S})$ is a stochastic function that maps a question $q$ to a distribution over prompt sequences $x$ (concretely: alternative rephrasings of the question), and $parse:\mathcal{S}\rightarrow\mathcal{A}$ is a deterministic function (e.g. regular expression matching) that maps a response sequence $y$ to an answer $a$ [^1] This parsing notation is very general. For instance, it can handle cases where the response contains chain-of-thought reasoning, in which case only the final answer is parsed. It also handles cases where the response fails or refuses to answer the question, provided the answer set $\mathcal{A}$ contains a corresponding ‘null’ answer. $\ell:\mathcal{A}\times\mathcal{A}\rightarrow\mathbb{R}_{\geq 0}$ is a non-negative loss function between the answer $a$ and the ground truth $a^{*}$ (lower is better). $\ell$ is readily linked to performance metrics like accuracy or ROUGE score.
@@
Our goal in this work is to develop an alternative method for selecting between responses from $phrase$ and $\pi$ to reduce the expected task loss beyond that achieved by NE and SC.
@@ 3 Representation Consistency
Raw model activations are known to be densely polysemantic anthropicsae , and human-understandable information can be encoded across model layers bereskamechanistic . This means that calculating similarity on them might include redundant, noisy information. To counter this potential issue, we propose an extra step for RC with Sparse-AutoEncoders (SAEs). They are neural network models operating on LLM activations to encode them into sparse and well-disentangled signals in SAEs’ latent space DBLP:conf/iclr/HubenCRES24 . Each layer of any LLM can be associated with an SAE trained on its activations. We denote the encoder of the SAE for the $l$-th layer as $f_{enc}^{l}:\mathbb{R}^{d_{\pi}}\rightarrow\mathbb{R}^{d_{SAE}}$, and ${d_{SAE}}$ is the number of latent dimensions of the SAE. Then, the only change from RC is the way consistency is calculated, with $sim$ function handling different input sizes:
@@ 4 Experiments
where $frequency$ is the same as that in RC (Equation 7). This method is inspired by how NLI (and embedding) models are used in uncertainty quantification DBLP:conf/iclr/KuhnGF23 ; AAAI25prompts and retrieval-augmented generations DBLP:conf/nips/rag . In our experiments, we use bge-m3-zeroshot-v2.0 chen2024bge ; laurer2023building for NLI with long sequence lengths. For all baselines, when there is a tie during answer aggregation, we select the last answer for reproducibility.
@@
We use both the raw model activations and the SAE-encoded activations to instantiate RC, respectively RC-D (dense) and RC-S (sparse). We follow the common practice of taking activations from the residual stream as in the mechanistic interpretability literature anthropicsae . For Gemma-2-9B-IT and Gemma-2-27B-IT models, we retrieve activations from the three layers on which GemmaScope SAEs are trained, located at 25%, 50%, 75% of model depth. For the other two models, their pretrained SAEs cover every layer, and we take activations from layers at 10%, 25%, 50%, 75%, and 90% of model depth. For each LLM generation, we cache the activation at the token location where the model is about to output the final answer choice, e.g., right after “the answer is: ”, and before “A”. In terms of the $sim$ function in Equations 6 and 10, we use cosine similarity to capture directions in the latent space. See Appendix A.2 for a detailed discussion on the implementation.
@@ 4.1 Task Performance Results
The task performances averaged for each model are summarised in Table 1. Figure 2 further reports detailed results for each model and dataset, averaged over prompt-sample configurations for 12 and 6 answers, respectively. See Appendix C.2 for full results on each configuration.
@@
Table 2 provides some insights into the optimal hyperparameters of RC-D and RC-S. We found that the most common optimal model layer for RC is at the middle, which coincides with recent research works stating that these layers often contain diverse high-level information templeton2024scaling ; DBLP:conf/emnlp/MenCJ00024 . $\lambda$ values are model- and dataset-specific, especially when the number of test samples is limited. Overall, the optimal relative importance of consistency and frequency in RC (Equation 5) is close to 1:1.
@@ 4.2 Answer Coherence Results
Next, we investigate the alignment between the consistency of model representations and what we, as humans, would perceive as coherent reasoning. We focus on the test cases (in each prompt-sampling configuration with 12 candidate responses) where the model produces 2 answers, and the number of responses supporting each answer is close, i.e., 6 vs. 6 and 5 vs. 7. To approximate our coherence notion, we employ LLM-as-a-judge. Specifically, we prompt the deepseek reasoner model (deepseek-R1) guo2025deepseek , asking which group of responses is more coherent in their reasoning. Using these results, we then record the percentage of times when the group of responses with better consistency in their representations (Equation 6, we use the 50% depth layer for every model) coincides with the coherence label. We perform the same evaluation with RC-E’s entailment probability (Equation 11) as a baseline.
@@
It is surprising to observe a substantial gap between RC-S and RC-D. This indicates that despite their similar task performances, they rely on different notions of consistency to achieve them. We can confirm that the model activations, after being processed by SAEs, align well with the common coherence notion. This is in line with the training objectives of SAEs - sparsify the dense signals in raw activations into sparse, potentially human-understandable concepts. We observe in our experiments that, among the many ($>10k$) SAE latent dimensions, only about 100 are activated (having non-zero values) at the target token position during generation. Therefore, a larger cosine similarity here means similar concepts are present when determining the final answer. This intuition naturally matches our notion of coherence. Similarly, RC-E’s entailment probability aims at obtaining the more coherent set, but is highly sensitive to the performance of the external NLI model on LLMs’ CoT-style responses.
@@ 5 Related Work
Test-time scaling refers to the problem of improving the task performance or output quality of LLMs after they have been trained zhang2025and . It is distinct from post-training DBLP:conf/nips/BrownMRSKDNSSAA20 , which updates model parameters via supervised fine-tuning bai2022training or reinforcement learning shao2024deepseekmath , instead seeking to make improvements without directly modifying the model. The prominent approaches are related to chain-of-thought reasoning, asking for some intermediate reasoning process DBLP:conf/nips/Wei0SBIXCLZ22 . This originally works on each response in a single-turn conversation. Further research extends this to ask for self-reflection and correction DBLP:conf/nips/MadaanTGHGW0DPY23 on the same LLM, forming multi-turn generations. There are also works to incorporate multiple LLMs performing similar conversations to reach more robust outputs, also in multiple turns khan2024debating ; DBLP:conf/icml/Du00TM24 . Works like self-consistency wang2023self , instead, obtain multiple responses from a single model (possibly in multi-turn taubenfeld2025confidence ) and then perform answer aggregation. Others bring a mix of the above approaches and intervene in the decoding process DBLP:conf/nips/YaoYZS00N23treeofthoughts ; DBLP:conf/aaai/BestaBKGPGGLNNH24graphofthoughts ; muennighoff2025s1 . Different to these approaches, we consider model internal activations for aggregating multiple LLM responses.
@@ 6 Conclusions
Exciting future works are envisaged following the introduction of RC. While we have investigated its integration with LLM response sets obtained by sampling and prompt rephrasing, the method could also be combined with more complex decoding-based scaling methods, e.g. tree-of-thoughts DBLP:conf/nips/YaoYZS00N23treeofthoughts . Consistency from multiple responses has been a useful signal for model training bai2022constitutional ; DBLP:conf/emnlp/DengWHWGSSXH22 ; consistencyrm , and using activations to aid training has been studied hao2024training ; shen2025codi . It would be interesting to explore how the consistency of activations can help in improving reasoning or model controllability. There are recent works separately using model internals DBLP:conf/iclr/insidellm or prompt rephrasings AAAI25prompts for uncertainty quantification and hallucination mitigation. Our method, combining both elements, could potentially be studied in this space too. Finally, our findings on the mismatch between raw model activation consistency and our understanding of coherence also highlight the need for further studies on the interpretability and transparency of LLMs hewitt2025we , possibly aided by SAEs.
@@ C.2 Task Performance Results per Configuration
Appendix D Impact Statements

### v2

@@ 2 Problem Definition
A QA task is a tuple $\tau=(\mathcal{Q},\mathcal{A},\rho,phrase,parse,\ell)$. $\mathcal{Q}$ is a set of questions and $\mathcal{A}$ is a set of answers. $\rho\in\Delta(\mathcal{Q}\times\mathcal{A})$ is a distribution from which questions $q$ and ground truth answers $a^{*}$ are sampled. $phrase:\mathcal{Q}\rightarrow\Delta(\mathcal{S})$ is a stochastic function that maps a question $q$ to a distribution over prompt sequences $x$ (concretely: alternative rephrasings of the question), and $parse:\mathcal{S}\rightarrow\mathcal{A}$ is a deterministic function (e.g. regular expression matching) that maps a response sequence $y$ to an answer aa. [^1] This parsing notation is very general. For instance, it can handle cases where the response contains chain-of-thought reasoning, in which case only the final answer is parsed. It also handles cases where the response fails or refuses to answer the question, provided the answer set $\mathcal{A}$ contains a corresponding ‘null’ answer. $\ell:\mathcal{A}\times\mathcal{A}\rightarrow\mathbb{R}_{\geq 0}$ is a non-negative loss function between the answer $a$ and the ground truth $a^{*}$ (lower is better). $\ell$ is readily linked to performance metrics like accuracy or ROUGE score.
@@
Our goal in this work is to develop an alternative method for selecting between responses from phrasephrase and $\pi$ to reduce the expected task loss beyond that achieved by NE and SC.
@@ 3 Representation Consistency
Raw model activations are known to be densely polysemantic anthropicsae , and human-understandable information can be encoded across model layers bereskamechanistic . This means that calculating similarity on them might include redundant, noisy information. To counter this potential issue, we propose an extra step for RC with Sparse-AutoEncoders (SAEs). They are neural network models operating on LLM activations to encode them into sparse and well-disentangled signals in SAEs’ latent space DBLP:conf/iclr/HubenCRES24 . Each layer of any LLM can be associated with an SAE trained on its activations. We denote the encoder of the SAE for the $l$-th layer as $f_{enc}^{l}:\mathbb{R}^{d_{\pi}}\rightarrow\mathbb{R}^{d_{SAE}}$, and $d_{SAE}$} is the number of latent dimensions of the SAE. Then, the only change from RC is the way consistency is calculated, with simsim function handling different input sizes:
@@
Incorporating SAEs only introduces slight additional computational costs to the standard RC. Apart from the memory space needed for caching the LLM activations, when $l$ and $\lambda$ are determined, the additional cost for an input text is one forward pass in an SAE. Specifically, the input to SAEs, an LLM’s residual activation at one token position, is of shape $(1,d_{\pi})$, and the number of parameters in each SAE is usually small (e.g., a pretrained SAE for the Gemma-2-9B model team2024gemma has 0.1B parameters gemmascope ). These indicate that the additional costs by SAEs are only a fraction of an LLM’s forward pass over a prompt.
@@ 4 Experiments
where frequencyfrequency is the same as that in RC (Equation 7). This method is inspired by how NLI (and embedding) models are used in uncertainty quantification DBLP:conf/iclr/KuhnGF23 ; AAAI25prompts and retrieval-augmented generations DBLP:conf/nips/rag . In our experiments, we use bge-m3-zeroshot-v2.0 chen2024bge ; laurer2023building for NLI with long sequence lengths. For all baselines, when there is a tie during answer aggregation, we select the last answer for reproducibility.
@@
We use both the raw model activations and the SAE-encoded activations to instantiate RC, respectively RC-D (dense) and RC-S (sparse). We follow the common practice of taking activations from the residual stream as in the mechanistic interpretability literature anthropicsae . For Gemma-2-9B-IT and Gemma-2-27B-IT models, we retrieve activations from the three layers on which GemmaScope SAEs are trained, located at 25%, 50%, 75% of model depth. For the other two models, their pretrained SAEs cover every layer, and we take activations from layers at 10%, 25%, 50%, 75%, and 90% of model depth. For each LLM generation, we cache the activation at the token location where the model is about to output the final answer choice, e.g., right after “the answer is: ”, and before “A”. In terms of the simsim function in Equations 6 and 10, we use cosine similarity to capture directions in the latent space. See Appendix A.2 for a detailed discussion on the implementation.
@@ 4.1 Task Performance Results
The task performances averaged for each model are summarised in Table 1. Figure 2 further reports detailed results for each model and dataset, averaged over prompt-sample configurations for 12 and 6 answers, respectively. Note that we include 4 models, 4 datasets, and 10 different prompt-sample configurations, making in total 160 sets of experiments. See Appendix C.2 for full results.
@@
Table 2 provides some insights into the optimal hyperparameters of RC-D and RC-S. We found that the most common optimal model layer for RC is at the middle, which coincides with recent research works stating that these layers often contain diverse high-level information templeton2024scaling ; DBLP:conf/emnlp/MenCJ00024 . $\lambda$ values are model- and dataset-specific, especially when the number of test samples is limited. Overall, the optimal relative importance of consistency and frequency in RC (Equation 5) is close to 1:1. See Appendix D for a detailed ablation study.
Table 3: Percentages where the correct/incorrect answer is associated with a higher consistency, as calculated by each method. The number after each model name is the number of interested cases matching the description.
Method/Model Gemma-2-2B- IT (497) Gemma-2-9B- IT (404) Gemma-2-27B- IT (113) Llama-3.1-8B- IT (195)
Embedding (baseline) 50.9%/49.1% 49.5%/50.5% 47.8%/52.2% 50.3%/49.7%
Dense (ours) 53.3%/46.7% 53.7%/46.3% 57.5%/42.5% 56.4%/43.6%
Sparse (ours) 53.9%/46.1% 54.7%/45.3% 58.4%/41.6% 56.4%/43.6%
We additionally perform an analysis validating the non-trivial usefulness of LLMs’ internal activations for aiding answer selection, compared against the surface-level textual outputs. For each model, we collected the test cases where among the 12 responses, 6 support one answer and the other 6 support another answer, and one of these two answers is the correct one. We count the number of times the correct vs. incorrect answer is associated with a higher consistency (dense activation consistency (ours, Equation 6), sparsified activation consistency (ours, Equation 10), and embedding consistency (the entailment probability baseline, Equation 11)). We aggregate the results for each model over all prompt-sample configurations and over the datasets. This is presented in Table 3. Note that the consistency measures are prone to the cases where the LLM is confidently wrong about some answer, so the absolute values of such measures could be tightly linked to the LLM’s performance. We observe that the correct answer constantly demonstrates a higher consistency calculated with internal activations, while this is not the case with textual embeddings. This could indicate that the activations help reveal information during the response generation process that is not included in the textual outputs.
@@ 4.2 Answer Coherence Results
Next, we investigate the alignment between the consistency of model representations and what we, as humans, would perceive as coherent reasoning. We focus on the test cases (in each prompt-sampling configuration with 12 candidate responses) where the model produces 2 answers, and the number of responses supporting each answer is close, i.e., 6 vs. 6 and 5 vs. 7. To approximate our coherence notion, we employ LLM-as-a-judge, and discuss LLMs’ suitability for this evaluation in Appendix E. Specifically, we prompt the deepseek reasoner model (deepseek-R1) guo2025deepseek , asking which group of responses is more coherent in their reasoning. Using these results, we then record the percentage of times when the group of responses with better consistency in their representations (Equation 6, we use the 50% depth layer for every model) coincides with the coherence label. We perform the same evaluation with RC-E’s entailment probability (Equation 11) as a baseline.
@@
It is surprising to observe a substantial gap between RC-S and RC-D. This indicates that despite their similar task performances, they rely on different notions of consistency to achieve them. We can confirm that the model activations, after being processed by SAEs, align well with the common coherence notion. This is in line with the training objectives of SAEs - sparsify the dense signals in raw activations into sparse, potentially human-understandable concepts. We observe in our experiments that, among the many (>10k>10k) SAE latent dimensions, only about 100 are activated (having non-zero values) at the target token position during generation. Therefore, a larger cosine similarity here means similar concepts are present when determining the final answer. This intuition naturally matches our notion of coherence. Similarly, RC-E’s entailment probability aims at obtaining the more coherent set, but is highly sensitive to the performance of the external NLI model on LLMs’ CoT-style responses.
@@ 5 Related Work
Test-time scaling refers to the problem of improving the task performance or output quality of LLMs after they have been trained zhang2025and . It is distinct from post-training DBLP:conf/nips/BrownMRSKDNSSAA20 , which updates model parameters via supervised fine-tuning bai2022training or reinforcement learning shao2024deepseekmath , instead seeking to make improvements without directly modifying the model. The prominent approaches are related to chain-of-thought reasoning, asking for some intermediate reasoning process DBLP:conf/nips/Wei0SBIXCLZ22 . This originally works on each response in a single-turn conversation. Further research extends this to ask for self-reflection and correction DBLP:conf/nips/MadaanTGHGW0DPY23 on the same LLM, forming multi-turn generations. There are also works to incorporate multiple LLMs performing similar conversations to reach more robust outputs DBLP:conf/emnlp/FreitagGF23 , with merged vocabularies across models DBLP:conf/nips/HuangFLXWL024 ; DBLP:conf/naacl/XuLZ24 , or in multiple turns khan2024debating ; DBLP:conf/icml/Du00TM24 . Works like self-consistency wang2023self , instead, obtain multiple responses from a single model (possibly in multi-turn taubenfeld2025confidence ) and then perform answer aggregation. Others bring a mix of the above approaches and intervene in the decoding process DBLP:conf/nips/YaoYZS00N23treeofthoughts ; DBLP:conf/aaai/BestaBKGPGGLNNH24graphofthoughts ; muennighoff2025s1 . Different to these approaches, we consider model internal activations for aggregating multiple LLM responses.
@@ 6 Conclusions
Exciting future works are envisaged following the introduction of RC. While we have investigated its integration with LLM response sets obtained by sampling and prompt rephrasing, the method could also be combined with more complex decoding-based scaling methods, e.g. tree-of-thoughts DBLP:conf/nips/YaoYZS00N23treeofthoughts . In this work, we examined RC on multiple-choice tasks. It would be an interesting direction to extend it to other reasoning tasks, possibly with more complex and open-ended generation forms. While designed for handling multiple responses from a single LLM, investigating RC’s transferability across multiple LLMs would also be desirable oozeeractivation . Consistency from multiple responses has been a useful signal for model training bai2022constitutional ; DBLP:conf/emnlp/DengWHWGSSXH22 ; consistencyrm , and using activations to aid training has been studied hao2024training ; shen2025codi . It would be interesting to explore how the consistency of activations can help in improving reasoning or model controllability. There are recent works separately using model internals DBLP:conf/iclr/insidellm or prompt rephrasings AAAI25prompts for uncertainty quantification and hallucination mitigation. Our method, combining both elements, could potentially be studied in this space too. Finally, our findings on the mismatch between raw model activation consistency and our understanding of coherence also highlight the need for further studies on the interpretability and transparency of LLMs hewitt2025we , possibly aided by SAEs.
@@ C.2 Task Performance Results per Configuration
Appendix D Ablation Analysis
We perform additional explorations on the RC method accuracy against the two hyperparameters in RC, namely $\lambda$ (ranging from -1 to 1), and LLM layer $l$ where the model activations are taken. While negative $\lambda$ values are not used in practice, we experiment with them to validate the usefulness of LLM activations. When $\lambda$ is negative, the evaluation function of RC (Equation 5) is calculated as: $\lambda\cdot consistency+(1-(abs(\lambda)))\cdot frequency$.
Figure 7: Ablation analysis on Llama3.1-8B-IT. The number following RC-D or RC-S indicates the index of layer $l$, see Table 5 for specific layer numbers.
Figure 8: Ablation analysis on Gemma-2-2B-IT. The number following RC-D or RC-S indicates the index of layer $l$, see Table 5 for specific layer numbers.
Figure 9: Ablation analysis on Gemma-2-9B-IT. The number following RC-D or RC-S indicates the index of layer $l$, see Table 5 for specific layer numbers.
Figure 10: Ablation analysis on Gemma-2-27B-IT. The number following RC-D or RC-S indicates the index of layer $l$, see Table 5 for specific layer numbers.
Figures 7 to 10 report the relative accuracy difference (vertical axis) against varying $\lambda$ values (horizontal axis) for each model and each dataset, averaged over four 6-response configurations. The majority vote result (the SC baseline) is obtained at $\lambda=0$. The only $\lambda$ value region with constant performance gain over the majority vote accuracy is when $0.25\leq\lambda\leq 0.85$ (roughly, depending on dataset and the choice of dense or sparse activations). The performance steadily drops with small fluctuations as $\lambda$ becomes more negative.
For every line, when $\lambda$ takes values between about [-0.25, 0.25], the accuracy stays constant (with very little fluctuations) at the SC result (majority vote, $\lambda=0$), because the frequency term is dominant in this region. When $\lambda$ increases from the above interval into the more positive region, the performance usually also starts to increase. The accuracy then quickly drops to a very low value because frequency no longer plays an important role (the method ends up choosing very infrequent answers). When $\lambda$ decreases from the above interval into the more negative region, in most cases, the performance will linearly decrease with fluctuations. At certain $\lambda$ locations, the performance can fluctuate to up to about 1% over the initial performance. RC-D reaches the performance peak and the performance drop at lower $\lambda$ values than RC-S.
The reason for a smaller performance drop at $\lambda=-1$ than when $\lambda=1$ is that the method tends to choose the majority answer in the most negative $\lambda$ region. This is because when the difference between the number of supporting responses for multiple answers is large (which is true for most test points), the majority answer tends to have a lower consistency due to the involvement of more responses, therefore is more likely to be selected by the RC method. For example, for a data point we have 12 responses, 3 are predicting answer A and 9 are predicting answer B, and answer B usually has a lower consistency because. It will more likely be selected by $\lambda=-1$ than by $\lambda=1$. This also highlights the importance of balancing consistency and frequency. We further note that very large absolute $\lambda$ values are impractical to use.
The differences between model layers $l$ are not particularly obvious, as similar performance gains can be observed for most configurations.
Appendix E LLMs’ Suitability to Evaluate Coherence
It should be noted that the results of Section 4.2 rely on the ability of LLMs to evaluate the concept of coherence. This reliance is supported by existing evidence: coherence is often used as a high-level evaluation metric during alignment training [70, 32], and LLMs are frequently employed as high-performing reward models [40]. To validate this capability in our specific context, we conducted a small-scale study. To validate their suitability in our context, we collected 10 examples from our experiments, each containing a question and 6 LLM responses evenly divided between two distinct answers (i.e., 3 responses per answer group). We manually labelled which group was more coherent and found that our LLM evaluator’s judgements perfectly matched our labels on these data points. We caution, however, that assessing coherence may become more difficult in practice with larger response sets. Below, we present an illustrative example from the MMLU dataset. In this case, the responses predicting answer B all exhibit identical reasoning and are therefore considered more coherent than the group predicting answer D.
Question: For their first exam, students in an AP Statistics class studied an average of 4 hours with a standard deviation of 1 hour. Almost everyone did poorly on the exam, and so for the second exam every student studied 10 hours. What is the correlation between the numbers of hours students studied for each exam?
Candidate answers: A: -1, B: 0, C: 0.4, D: 1
Correct answer: B: 0
Group 1 responses predicting D:1
  1. 1.
    Here’s how to think about this: **Correlation measures the strength and direction of a linear relationship between two variables.** **In this case, the two variables are the number of hours studied for the first exam and the number of hours studied for the second exam.** **Notice that every student studied 10 hours for the second exam, regardless of how much they studied for the first exam.** This means there’s a perfect, direct relationship: the number of hours studied for the second exam is always 10, no matter what the first exam study time was. **A perfect, direct relationship has a correlation of 1.** [The answer is: D]
  2. 2.
    Here’s how to determine the correlation: **Understanding Correlation:** Correlation measures the strength and direction of the linear relationship between two variables. A correlation of -1 indicates a perfect negative linear relationship, 0 indicates no linear relationship, and 1 indicates a perfect positive linear relationship. **Analyzing the Scenario:** In this case, we’re looking at the relationship between hours studied for the first exam and hours studied for the second exam. **Key Observation:** Every student increased their study time by a fixed amount (9 hours) for the second exam. This means there’s a perfect, direct relationship between the two study times. If a student studied 3 hours for the first exam, they studied 12 hours for the second. If they studied 5 hours for the first, they studied 14 hours for the second, and so on. **Conclusion:** Since there’s a perfect, direct relationship, the correlation is 1. [The answer is: D]
  3. 3.
    Here’s how to determine the correlation: **Understanding Correlation:** Correlation measures the strength and direction of the linear relationship between two variables. A correlation of -1 indicates a perfect negative linear relationship, 0 indicates no linear relationship, and 1 indicates a perfect positive linear relationship. **Analyzing the Situation:** In this case, every student increased their study time by a fixed amount (from an average of 4 hours to 10 hours). This means the relationship between study time for the first exam and the second exam is perfectly linear. **Perfect Linear Relationship:** Since the study time for the second exam is directly determined by the study time for the first exam (plus 6 hours), the correlation is 1. [The answer is: D]
Group 2 responses predicting B:0
  1. 1.
    Here’s how to determine the correlation: **Understanding Correlation:** Correlation measures the strength and direction of the linear relationship between two variables. A correlation of -1 indicates a perfect negative linear relationship, 0 indicates no linear relationship, and 1 indicates a perfect positive linear relationship. **Analyzing the Situation:** In this case, we’re looking at the number of hours studied for the first exam and the number of hours studied for the second exam. Since *all* students studied 10 hours for the second exam, regardless of how much they studied for the first exam, there is no linear relationship between the two. **The Answer:** [The answer is: B]
  2. 2.
    Here’s how to reason through this: **Correlation measures the strength and direction of a linear relationship between two variables.** **In this case, we’re looking at the relationship between hours studied for the first exam and hours studied for the second exam.** **The key information is that every student studied 10 hours for the second exam, regardless of how much they studied for the first.** This means there’s no variation in study time for the second exam based on the first exam’s study time. **Since there’s no linear relationship (hours studied for the first exam doesn’t predict hours studied for the second), the correlation is 0.** [The answer is: B]
  3. 3.
    Here’s how to determine the correlation: **Understanding Correlation:** Correlation measures the strength and direction of a linear relationship between two variables. A correlation of -1 indicates a perfect negative linear relationship, 0 indicates no linear relationship, and 1 indicates a perfect positive linear relationship. **Analyzing the Situation:** In this case, we know that everyone increased their study time for the second exam. Since every student studied 10 hours for the second exam, regardless of how much they studied for the first exam, there is no linear relationship between the two. [The answer is: B]
Appendix F Impact Statements